OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = 3(1.02)x x g(x) -1 -4 0 6 1 8 2 10 Which statement best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus y-intercept of f(x).

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

ok. First draw a table and so that the numbers starts talking to you :)

OpenStudy (anonymous):

I drew it on my paper, can you walk me threw it?

OpenStudy (anonymous):

yah :)

OpenStudy (anonymous):

Wait I'll draw one here

OpenStudy (anonymous):



OpenStudy (anonymous):

now try to find the relation between x and g(x)

OpenStudy (anonymous):

It's a little difficult, but we can do it, right? Let's try.

OpenStudy (anonymous):

I have to go very soon, just to let you know :) I have a few more questions that are easier. :)

OpenStudy (anonymous):

ok let's try this just one more time. And if we don't get it, we will move on to the next

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

While we wait for him, lets go to the next one?

OpenStudy (anonymous):

Bill borrowed the same amount of money from Linda and Drake. The table below shows the amount, in dollars, that Bill would owe them after different numbers of years: Year 1 2 3 4 Linda 207 214 221 228 Drake 206 212.18 218.55 225.10 Which statement is true about the money Bill would owe Linda and Drake after 30 years? He would owe Linda twice the amount he borrowed. He would owe both the same amount of money. He would owe Linda more money. He would owe Drake more money.

OpenStudy (anonymous):

See how x is related to f(x) and g(x)

OpenStudy (anonymous):

f(x) is Linda's money and g(x) is Drake's money

OpenStudy (anonymous):

On Linda's side, you can see that there is an increase of 7 dollars per years

OpenStudy (anonymous):

*year

OpenStudy (anonymous):

So C? she owes more?

OpenStudy (anonymous):

Yes. Because the increase in her amount is more than the increase in Drake's amount. Right :)

OpenStudy (anonymous):

The graph of f(x) = (0.5)x is replaced by the graph of g(x) = (0.5)x - k. If g(x) is obtained by shifting f(x) down by 3 units, the value of k is ________. Numerical Answers Expected!

OpenStudy (anonymous):

Im so close to being done!!!! lol

OpenStudy (anonymous):

f(x) = (0.5)x g(x) = (0.5) - k If the graph shifts down by 3 units, k should be 3 units,.

OpenStudy (anonymous):

because from the data given, g(x) = f(x) - 3 = (0.5)x - 3

OpenStudy (anonymous):

So answer is 3?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

Heidi is preparing for the national gymnastics competition. The table below shows the number of hours Heidi spent preparing for a gymnatics competition over a period of five months: Month 1 2 3 4 5 Hours 2 3.5 5 6.5 8 Did Heidi increase the number of hours of practice linearly or exponentially? Linearly, because the table shows an equal increase in hours for an equal increase in months Exponentially, because the table shows an equal increase in hours for an equal increase in months Linearly, because the table shows that hours increase by an equal factor for an equal increase in months Exponentially, because the table shows that hours increase by an equal factor for an equal increase in months

OpenStudy (anonymous):

Her no. of working hours per month is 0.5 + 1.5x

OpenStudy (anonymous):

There are no exponents involved. Just linear addition depending on the number of months

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

So the answer is a

OpenStudy (anonymous):

Oh.. I see. xD

OpenStudy (anonymous):

It can't be b because there are no exponents in the solution. You are not multiplying them. Just adding the numbers. So it will not be exponential.

OpenStudy (anonymous):

The tenth, eleventh, and twelfth terms of a sequence are shown in the table below: Term number 10 11 12 Term 42 47 52 Which of the following shows the first five terms of the sequence? 3, -2, -7, -12, -17 2, 7, 12, 17, 22 -3, 2, 7, 12, 17 -2, -7, -12, -17, -22

OpenStudy (anonymous):

There is an increase of 5 for every succeeding term

OpenStudy (anonymous):

B

OpenStudy (anonymous):

If the 10th term is 42, first term is 42 - 9(x) where x is the increase in the value of terms right?

OpenStudy (anonymous):

Wht?

OpenStudy (anonymous):

Ok. Let the first term be A

OpenStudy (anonymous):

And the increase is 5 units every time. Ok?

OpenStudy (anonymous):

ALright

OpenStudy (anonymous):

so.....its...

OpenStudy (anonymous):

So 2nd term will be A + 5 right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

3rd term is A + 10

OpenStudy (anonymous):

If you keep on doing this, you will get that 10th term is A + 45

OpenStudy (anonymous):

So A + 45 = 42

OpenStudy (anonymous):

Wait whats the answer?

OpenStudy (anonymous):

So A = 42-45 = -3. If the first term is -3, 2nd term is 2, 3rd is 7 and so on

OpenStudy (anonymous):

So the answer is c

OpenStudy (anonymous):

Jess plans to increase the number of hours she spends doing exercises. She can increase the hours in the following ways: Month 1 2 3 4 5 Option 1 0.5 0.8 1.1 1.4 1.7 Option 2 0.5 1 2.0 4.0 8.0 Option 3 0.5 0.55 0.61 0.67 0.74 If Jess wants to increase the hours linearly, which option(s) should she choose? Only option 1, because it shows equal increases in equal intervals of time Only option 2, because it shows equal increases in equal intervals of time Either option 1 or option 2, because they show increases in time by the same percentage Either option 2 or option 3, because they show increases in time by the same percentage

OpenStudy (anonymous):

This one and more more!

OpenStudy (anonymous):

This one, and one more*!

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Where is the increase uniform?

OpenStudy (anonymous):

In option 1 right?

OpenStudy (anonymous):

it's increasing by 0.3 in each case

OpenStudy (anonymous):

Red u there?

OpenStudy (anonymous):

Yes sorry im here

OpenStudy (anonymous):

In option 2, it is increasing by 2, 4, 8 TIMES, which is exponential. In option 3, the increase is not uniform. So it's not linear. So the answer is option a, option 1

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 8 4x −5y = 2 3x − 8y = 4 4x − 5y = 2 13x − 8y = 26 4x −5y = 2 10x + 3y = 15 Which pair of systems will have the same solution? System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2 System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1

OpenStudy (anonymous):

Did u understand the previous answer?

OpenStudy (anonymous):

Yes I did

OpenStudy (anonymous):

hey brb

OpenStudy (anonymous):

wait!!!! I have one more!

OpenStudy (anonymous):

just one minute i'll be back after that

OpenStudy (anonymous):

Alright